# Efficient Deep Spiking Multi-Layer Perceptrons with Multiplication-Free Inference

### v1

@@ Abstract
Advancements in adapting deep convolution architectures for Spiking Neural Networks (SNNs) have significantly enhanced image classification performance and reduced computational burdens. However, the inability of Multiplication-Free Inference (MFI) to harmonize with attention and transformer mechanisms, which are critical to superior performance on high-resolution vision tasks, imposes limitations on these gains. To address this, our research explores a new pathway, drawing inspiration from the progress made in Multi-Layer Perceptrons (MLPs). We propose an innovative spiking MLP architecture that uses batch normalization to retain MFI compatibility and introduces a spiking patch encoding layer to reinforce local feature extraction capabilities. As a result, we establish an efficient multi-stage spiking MLP network that effectively blends global receptive fields with local feature extraction for comprehensive spike-based computation. Without relying on pre-training or sophisticated SNN training techniques, our network secures a top-1 accuracy of 66.39% on the ImageNet-1K dataset, surpassing the directly trained spiking ResNet-34 by 2.67%. Furthermore, we curtail computational costs, model capacity, and simulation steps. An expanded version of our network challenges the performance of the spiking VGG-16 network with a 71.64% top-1 accuracy, all while operating with a model capacity 2.1 times smaller. Our findings accentuate the potential of our deep SNN architecture in seamlessly integrating global and local learning abilities. Interestingly, the trained receptive field in our network mirrors the activity patterns of cortical cells.
@@ I Introduction
Over the years, Convolutional Neural Networks (CNNs) have garnered considerable success within the field of computer vision. However, the advent and subsequent accomplishments of Transformer [1] in natural language processing have led to the emergence of the Vision Transformer (ViT) [2] as a promising contender. The ViT model replaces the convolution operation in CNNs with the self-attention operation from the Transformer, enabling it to model visual relationships across different spatial locations of an image. ViT and its subsequent developments [3, 4] have showcased performance on par with, or even superior to, CNNs. Unlike CNNs, which necessitate intricate convolution kernel designs, ViTs utilize several standard Transformer blocks, thus minimizing hand-crafted manipulations and reducing inductive biases.
More recently, the MLP-Mixer [5] research proposed a simpler approach, leveraging Multilayer Perceptrons (MLP) to further mitigate induction bias. The fundamental block of MLP-Mixer consists of two components: the channel-mixing MLP and the token-mixing MLP. The channel-mixing module performs MLP calculations on the channel dimension of the feature map, facilitating information exchange across different channels. Simultaneously, the token-mixing module conducts MLP calculations across varying positions of the feature map, enabling communication over the spatial dimension. By harmonizing these two modules, the MLP-Mixer can efficiently extract and integrate features from both dimensions, enhancing the model’s overall performance. This methodology presents a more straightforward route to reduce induction bias in comparison with previous models.
From a communication perspective, Spiking Neural Networks (SNNs) exhibit a finely-tuned mechanism for managing the exchange of information. Rooted in computational neuroscience, SNNs have been extensively utilized for the modeling of brain dynamics and functions [6, 7, 8, 9, 10]. Recently, the successes and challenges experienced by deep Artificial Neural Networks (ANNs) in tackling machine learning problems have fostered a growing curiosity towards SNNs. Researchers are exploring them as potential alternatives and are keen to leverage their bio-inspired properties to address similar issues.
The competence of SNNs has been convincingly demonstrated, with their performance proving to be on par with their artificial counterparts [11, 12, 13, 14, 15, 16, 17, 18, 19]. This has been achieved through the astute application of adaptive learning algorithms [20, 21, 22, 23] and the integration of efficient architectures borrowed from ANNs. These architectures span from the Boltzmann machines [24], prominent during the nascent stages of deep learning, to the currently dominant CNNs [25]. Furthermore, to identify task-specific cells or network structures for SNNs, recent research has employed Neural Architecture Search (NAS) [26, 27]. This approach not only strengthens the performance of SNNs but also broadens their potential application scope within the realm of machine learning.
@@
The rest of this paper is organized as follows: Section II delves into the related works MLPs and SNNs, two distinct deep learning approaches; Section III first provides an overview of our network architecture, followed by a detailed discussion on the vital Spiking MLP Block module, including the LIF neurons utilized therein; Section IV describes the tests conducted on three distinct datasets (ImageNet-1K, CIFAR10/100, and CIFAR10-DVS) for classification tasks, and presents our findings, including results from ablation, network sparsity, and neuronal weight visualization experiments; finally, Section V encapsulates the results and explores the implications of our work.
@@ II-A Multi-Layer Perceptrons in Deep Learning
In the quest for an alternative architecture, the ground-breaking work of the MLP-Mixer [5] offered a model relying solely on Multi-Layer Perceptrons (MLPs), eschewing both convolution and self-attention layers. The MLP-Mixer is comprised of two unique parts - the channel-mixing MLP and the token-mixing MLP. The channel mixing module primarily focuses on conducting MLP computations within the channel dimension of the feature map, leading to an exchange of information among various channels and thereby fostering the extraction of effective features. Conversely, the token mixing module carries out MLP computations among different feature map positions. This enhances communication across spatial dimensions, fostering integration of diverse features from various map locations. Together, these modules equip the MLP-Mixer model with a robust proficiency for feature extraction and integration from both dimensions.
@@ II-B Spiking Neural Networks in Deep Learning
Spiking Neural Networks (SNNs) have established their prowess in effective information processing, primarily due to their bio-inspired mechanisms. Deriving their roots from computational neuroscience, SNNs are frequently employed to model brain dynamics and functions [6, 7, 8, 9, 10]. As machine learning evolves through its successes and challenges, SNNs have emerged as promising alternatives with researchers exploring their biological properties for functional advantages in similar applications.
Although SNNs were not originally designed for gradient-based supervised learning, their bio-inspired architecture and capability to process spiking neural activity have led to a surge in popularity. Traditional training methods for SNNs, such as Spike Timing-Dependent Plasticity (STDP) [40], while effective, pose limitations in incorporating global information. This shortfall can slow convergence speed and impede applicability to large models. Further, STDP can result in unpredictable neural behavior, complicating the understanding and optimization of network performance.
@@ III-A Network Architecture
Figure 1: The overall network architecture. The multi-stage network is downsampled with a spiking patch encoding (SPE) module at each stage. Within each stage, the SPE is followed by a sequence of spiking MLP-Mixers with identical architecture, each containing a spiking token block with axial sampling and a spiking channel block with full sampling.
(a) Spiking Token Block
(b) Spiking Channel Block
(c) Spiking MLP Block
@@
Specifically, our approach starts with an RGB input image of size 3×3$\times$3 ×H×$\times$×W. The original MLP-Mixer breaks this image down into non-overlapping patches of size $p\times p$, subsequently projecting the new channel dimension of $3\times p^{2}$ to a hidden dimension of $C_{1}$. This process parallels a 2D convolution operation with a kernel and stride size of $p$, and an output dimension of $C_{1}$. Although the MLP-Mixer benefits from a global receptive field, it lacks the necessary inductive bias for learning local features. To address this deficiency, we develop a spiking patch encoding (SPE) module to replace the original patch partitioning approach. We construct the SPE with a spike-based directed acyclic graph, following the design outlined in a recent work [18], as illustrated in Figure 3.
@@ III-B Spiking MLP-Mixer
A straightforward adoption of existing MLP block designs from artificial neural networks (ANNs) by substituting the real-valued activation function with a spiking activation function (SAF) results in a violation of the MFI principle of spike-based computation. In both the token and channel blocks of the original MLP-Mixer, the fully connected (FC) operation is carried out on the layer normalization (LN) state of the feature map, leading to real-valued matrix multiplication. To rectify this, we move the normalization process to after the FC operation and replace LN with batch normalization (BN). This adjustment enables the parameters of the latter to be integrated with the linear projection weights during inference, aligning with the MFI principle [52, 42].
The spiking MLP-Mixer comprises a token block and a channel block, each of which includes multiple FC layers with spiking activation, as depicted in Fig. 1(b) and Fig. 1(c).
@@ III-C LIF Spiking Neuron
[…]
@@ IV Experiments
In order to thoroughly evaluate the performance and capabilities of our model, a variety of image classification benchmarks are utilized. These include the ImageNet-1K [58], CIFAR10/100 [59], and CIFAR10-DVS [60] datasets. The SpikingJelly [61] toolkit, complemented by its cupy backend, enables rapid simulation of Leaky Integrate-and-Fire (LIF) neurons. The Spiking Pooling Equilibrium (SPE) module incorporated in our architecture utilizes $3\times 3$ convolution operations. The stride for input-to-node connections is set to 2, while node-to-node connections have a stride of 1. To avoid statistical bias, the batch mean and variance are synchronized across each device, following the approach suggested by [15].
The following sections will present results from the aforementioned datasets, discuss an ablation study, assess the network sparsity and computational cost, and finally visualize our findings.
@@ IV-A Results on ImageNet-1K
The expansion ratio of the channel FC layer is set to $\alpha=3$. To study the influence of the SPE module, we also train an MLP-S model with the original patch partition approach while other parts of the network remaining the same. The simulation step of SNN is set to $T=4$. To better compare with other SNNs with larger time steps, we use time inheritance training (TIT) [17] and obtain the $T=6$ result for MLP-SPE-T, where we initialize the network with the pre-trained MLP-SPE-T ($T=4$) model and fine-tune for 50 epochs with $T=6$ using a cosine learning rate decaying from 0.1 to 0.
@@
Table I: Results on ImageNet-1K. T denotes simulation length.
@@
It can be observed that ANN-SNN conversion methods can indeed achieve high accuracy, but they require exceedingly lengthy simulation steps. Hybrid training methods manage to reduce the simulation steps but remain considerably longer than methods using direct training. Among directly trained methods, and considering similar network capacity, MLP-SPE-T significantly outperforms STBP-tdBN ResNet-34 by 2.67%, all the while requiring fewer simulation time steps. Post-TIT, the model achieves an accuracy of 69.09% with $T=6$, exceeding ResNet-34-large and VGG-16 models by 2.04% and 0.09% respectively, while maintaining a model capacity more than 3 and 5.5 times smaller. These results underscore the superior efficiency of our architecture.
The enhancement observed in MLP-SPE-S over MLP-S confirms that the SPE module can substantially improve network performance. This result attests to the efficacy of amalgamating global receptive fields with local feature extraction. However, it is important to note that our method is more accurately compared with the STBP-tdBN method since both use the same conventional temporal averaged loss function. The Temporal Error Transport (TET) method, which employs a temporal moment-wise loss function, has been demonstrated to outperform the temporal averaged method [17].
@@ IV-B Results on CIFAR
Our results are benchmarked against other state-of-the-art SNNs, as depicted in Table II. Despite a larger network size, our model sets new records on both datasets, outperforming existing SNNs significantly. While comparing our results with directly trained SNNs might seem somewhat disproportionate, the outcomes convincingly illustrate the efficacy of the spiking MLP network as a pre-trained model applicable to other datasets.
Table II: Results on CIFAR. T denotes simulation length.
@@
CIFAR10 Diet-SNN [30] ResNet-20 10 $92.54$
@@
CIFAR100 Diet-SNN [30] ResNet-20 5 $64.07$
@@ IV-C Results on CIFAR10-DVS
Neuromorphic datasets generally possess higher noise levels in comparison to static datasets, which heightens the risk of overfitting in well-optimized SNNs. Among all mainstream neuromorphic datasets, CIFAR10-DVS is recognized as one of the most challenging, presenting approximately 0.9k training samples for each label. Recent literature suggests an inclination towards complex architectures to handle this dataset effectively, albeit this strategy carries an inherent risk of overfitting without correspondingly enhancing the model’s accuracy. Therefore, a pressing need exists for novel methodologies that can abate the influence of noise and tackle the difficulties inherent in working with CIFAR10-DVS.
As with the CIFAR10/100 datasets, we employ the spiking MLP-SPE-T network, pre-trained on the ImageNet-1K dataset, as the pre-trained model for the CIFAR10-DVS dataset. We reset the final classification output layer of the network, and subsequently fine-tune it on the resized $224\times 224$ CIFAR10-DVS dataset. Our fine-tuning strategy employs a cosine decay learning rate, starting with an initial value of 0.1 and progressively reducing to zero over 100 epochs. SGD, with a momentum of 0.9, is utilized as the optimizer.
@@
Table III: Comparison on CIFAR10-DVS. T denotes simulation length.

### v3

@@ Abstract
Advancements in adapting deep convolution architectures for Spiking Neural Networks (SNNs) have significantly enhanced image classification performance and reduced computational burdens. However, the inability of Multiplication-Free Inference (MFI) to align with attention and transformer mechanisms, which are critical to superior performance on high-resolution vision tasks, imposing limitations on these gains. To address this, our research explores a new pathway, drawing inspiration from the progress made in Multi-Layer Perceptrons (MLPs). We propose an innovative spiking MLP architecture that uses batch normalization to retain MFI compatibility and introducing a spiking patch encoding layer to enhance local feature extraction capabilities. As a result, we establish an efficient multi-stage spiking MLP network that blends effectively global receptive fields with local feature extraction for comprehensive spike-based computation. Without relying on pre-training or sophisticated SNN training techniques, our network secures a top-1 accuracy of 66.39% on the ImageNet-1K dataset, surpassing the directly trained spiking ResNet-34 by 2.67%. Furthermore, we curtail computational costs, model parameters, and simulation steps. An expanded version of our network compares with the performance of the spiking VGG-16 network with a 71.64% top-1 accuracy, all while operating with a model capacity 2.1 times smaller. Our findings highlight the potential of our deep SNN architecture in effectively integrating global and local learning abilities. Interestingly, the trained receptive field in our network mirrors the activity patterns of cortical cells. Source codes are publicly accessible at https://github.com/EMI-Group/mixer-snn.
@@ I Introduction
Over the years, Convolutional Neural Networks (CNNs) have garnered considerable success within the field of computer vision. However, the advent and subsequent accomplishments of Transformer [1] in natural language processing have led to the emergence of the Vision Transformer (ViT) [2] as a promising contender. The ViT model replaces the convolution operation in CNNs with the self-attention operation from the Transformer, enabling it to model visual relationships across different spatial locations of an image. ViT and its later versions [3, 4] have showcased performance on par with, or even superior to, CNNs. Unlike CNNs, which require intricate convolution kernel designs, ViTs utilize several standard Transformer blocks, thus minimizing manual manipulations and reduce inductive biases.
More recently, the MLP-Mixer [5] researchers proposed a simpler approach, leveraging Multilayer Perceptrons (MLP) to further mitigate induction bias. The fundamental block of MLP-Mixer consists of two components: the channel-mixing MLP and the token-mixing MLP. The channel-mixing module performs MLP calculations on the channel dimension of the feature map, facilitating information exchange across different channels. Simultaneously, the token-mixing module conducts MLP calculations across varying positions of the feature map, enabling communication over the spatial dimension. By harmonizing these two modules, the MLP-Mixer can efficiently extract and integrate features from both dimensions, enhancing the model’s overall performance. This methodology presents a simpler approach to reduce induction bias in comparison with previous models.
Rooted in computational neuroscience, SNNs have been extensively utilized for the modeling of brain dynamics and functions [6, 7, 8, 9, 10]. Recently, the successes and challenges experienced by deep Artificial Neural Networks (ANNs) in tackling machine learning problems have fostered a growing curiosity towards SNNs. Researchers are exploring them as alternatives and are keen to leverage their bio-inspired properties to address similar issues.
The competence of SNNs has been convincingly demonstrated, with their performance showing promising results [11, 12, 13, 14, 15, 16, 17, 18, 19]. This has been achieved through the astute application of adaptive learning algorithms [20, 21, 22, 23] and the integration of efficient architectures borrowed from ANNs. These architectures span from the Boltzmann machines [24], prominent during the nascent stages of deep learning, to the currently dominant CNNs [25]. Furthermore, to identify task-specific cells or network structures for SNNs, recent research has employed Neural Architecture Search (NAS) [26, 27]. This approach not only strengthens the performance of SNNs but also broadens their potential application scope within the realm of machine learning.
@@
The rest of this paper is organized as follows: Section II delves into the related works MLPs and SNNs, two distinct deep learning approaches; Section III first provides an overview of our network architecture, followed by a detailed discussion on the vital Spiking MLP Block module; Section IV describes the tests conducted on three distinct datasets (ImageNet-1K, CIFAR10/100, and CIFAR10-DVS) for classification tasks, and presents our findings, including results from ablation, network spiking rate, and neuronal weight visualization experiments; finally, Section V encapsulates the results and explores the implications of our work.
@@ II-A Multi-Layer Perceptrons in Deep Learning
In the quest for an alternative architecture, the ground-breaking work of the MLP-Mixer [5] offered a model relying solely on MPLs, eschewing both convolution and self-attention layers. The MLP-Mixer consists of two unique parts - the channel-mixing MLP and the token-mixing MLP. The channel mixing module primarily focuses on conducting MLP computations within the channel dimension of the feature map, leading to an exchange of information among various channels and thereby fostering the extraction of effective features. Conversely, the token mixing module carries out MLP computations among different feature map positions. This enhances communication across spatial dimensions, fostering integration of diverse features from various map locations. Together, these modules equip the MLP-Mixer model with a robust proficiency for feature extraction and integration from both dimensions.
@@ II-B Spiking Neural Networks in Deep Learning
Spiking Neural Networks (SNNs) have established their prowess in effective information processing, primarily due to their bio-inspired mechanisms [40, 41, 42, 43]. Deriving their roots from computational neuroscience, SNNs are frequently employed to model brain dynamics and functions [6, 7, 8, 9, 10]. As machine learning evolves through its successes and challenges, SNNs have emerged as promising alternatives with researchers exploring their biological properties for functional advantages in similar applications.
Although SNNs were not originally designed for gradient-based supervised learning, their bio-inspired architecture and capability to process spiking neural activity have led to a surge in popularity. Traditional training methods for SNNs, such as Spike Timing-Dependent Plasticity (STDP) [44], while effective, pose limitations in incorporating global information. This shortfall can slow convergence speed and impede applicability to large models.
@@ III-A Network Architecture
Figure 1: The overall network architecture. The multi-stage network is downsampled with an SPE module at each stage. Within each stage, the SPE is followed by a sequence of spiking MLP-Mixers with identical architecture, each containing a spiking token block with axial sampling and a spiking channel block with full sampling.
(a) Spiking Token Block.
(b) Spiking Channel Block.
(c) Spiking MLP Block.
@@
Specifically, our approach starts with an RGB input image of size 3×3$\times$3 ×H×$\times$×W. The original MLP-Mixer breaks this image down into non-overlapping patches of size $p\times p$, subsequently projecting the new channel dimension of $3\times p^{2}$ to a hidden dimension of $C_{1}$. This process parallels a 2D convolution operation with a kernel and stride size of $p$, and an output dimension of $C_{1}$. Although the MLP-Mixer benefits from a global receptive field, it lacks the necessary inductive bias for learning local features. To address this deficiency, we develop an SPE module to replace the original patch partitioning approach. We construct the SPE with a spike-based directed acyclic graph, following the design outlined in a recent work [18], as illustrated in Figure 3.
@@ III-B Spiking MLP-Mixer
A straightforward adoption of existing MLP block designs from artificial neural networks (ANNs) by substituting the real-valued activation function with a spiking activation function (SAF) results in a violation of the MFI principle of spike-based computation. In both the token and channel blocks of the original MLP-Mixer, the fully connected (FC) operation is carried out on the layer normalization (LN) [56] state of the feature map, leading to real-valued matrix multiplication. To rectify this, we move the normalization process to after the FC operation and replace LN with batch normalization (BN). This adjustment enables the parameters of the latter to be integrated with the linear projection weights during inference, aligning with the MFI principle [57, 46].
The spiking MLP-Mixer comprises a token block and a channel block, each of which includes multiple FC layers with spiking activation, as depicted in Fig. 2.
@@ III-C LIF Spiking Neuron
where $t$ signifies the time step, $\tau$ represents the membrane time constant, and ${\bm{u}}$ is the membrane potential, denoted by a bold, italicized letter to represent a vector. ${\bm{y}}$ stands for the spike output and $g$ is a threshold function. ${\bm{z}}^{t}={\bm{W}}{\bm{y}}^{t,\mathrm{pre}}$ refers to the synaptic input, where ${\bm{W}}$ is the weight matrix, and ${\bm{y}}^{t,\mathrm{pre}}$ signifies the afferent spikes originating from pre-synaptic neurons.
@@ IV Experiments
In order to thoroughly evaluate the performance and capabilities of our model, a variety of image classification benchmarks are utilized. These include the ImageNet-1K [63], CIFAR10/100 [64], and CIFAR10-DVS [65] datasets. The SpikingJelly [66] toolkit, complemented by its cupy backend, enables rapid simulation of Leaky Integrate-and-Fire (LIF) neurons. The SPE module incorporated in our architecture utilizes $3\times 3$ convolution operations. The stride for input-to-node connections is set to 2, while node-to-node connections have a stride of 1. To avoid statistical bias, the batch mean and variance are synchronized across each device, following the approach suggested by [15].
The following sections will present results from the aforementioned datasets, discuss an ablation study, assess the network spiking rate and computational cost, and finally visualize our findings.
@@ IV-A Results on ImageNet-1K
The expansion ratio of the channel FC layer is set to $\alpha=3$. To study the influence of the SPE module, we also train an MLP-S model with the original patch partition approach while other parts of the network remaining the same. The simulation step of SNN is set to $T=4$, where $T$ denotes the simulation step. To better compare with other SNNs with larger time steps, we use time inheritance training (TIT) [17] and obtain the $T=6$ result for MLP-SPE-T, where we initialize the network with the pre-trained MLP-SPE-T ($T=4$) model and fine-tune for 50 epochs with $T=6$ using a cosine learning rate decaying from 0.1 to 0.
@@
Table I: Results on ImageNet-1K (T denotes simulation length)
@@
SpikFormer (SNN) [72] Spikformer-8-512 30M 4 $73.38$
SparseMLP (ANN) [36] sMLPNet-T 24M - $81.9$
It can be observed that ANN-SNN conversion methods can indeed achieve high accuracy, but they require exceedingly lengthy simulation steps. Hybrid training methods manage to reduce the simulation steps but remain considerably longer than methods using direct training. Among directly trained methods, and considering similar network capacity, MLP-SPE-T significantly outperforms STBP-tdBN ResNet-34 by 2.67%, all the while requiring fewer simulation time steps. Post-TIT, the model achieves an accuracy of 69.09% with $T=6$, exceeding ResNet-34-large, VGG-16 and the DARTs-based SpikeDHS models by 2.04%, 0.09%, 1.13% respectively, while maintaining a model capacity more than 3, 5.5 and 2.3 times smaller. These results underscore the superior efficiency of our architecture.
The enhancement observed in MLP-SPE-S over MLP-S confirms that the SPE module can substantially improve network performance. This result attests to the efficacy of amalgamating global receptive fields with local feature extraction. However, it is important to note that our method is more accurately compared with the STBP-tdBN method since both use the same conventional temporal averaged loss function. The Temporal Error Transport (TET) method, which employs a temporal moment-wise loss function, has been demonstrated to outperform the temporal averaged method [17]. Note that the work of [31] has applied temporal self-recurrent non-spiking neurons in MLP architecture, achieving an accuracy similar to transformer-based models on ImageNet-1K. However, due to the use of floating-point valued neurons, the model has no significant advantage in terms of energy cost compared with other ANNs. The SparseMLP [36] in the ANN domain and SpikFormer [72] using attention mechanism have significant higher accuracy compared with our model under similar parameter number. However, the sparse spiking activation of our architecture renders much lower energy cost, as demonstrated in Table VI.
@@ IV-B Results on CIFAR
Our results are benchmarked against other state-of-the-art SNNs, as depicted in Table II. Despite a larger network size, our model sets new records on both datasets, outperforming existing SNNs significantly. While comparing our results with directly trained SNNs might seem somewhat disproportionate, the outcomes convincingly illustrate the efficacy of the spiking MLP network as a pre-trained model applicable to other datasets. Note that our model also outperforms [26], a light weight SNN specifically optimized for balanced accuracy and energy cost, under much smaller timestep.
Table II: Results on CIFAR (T denotes simulation length)
@@
CIFAR10 TerMapping [73] VGG 2800 $93.75$
AugMapping [73] VGG 300 $93.90$
S-ResNet [69] ResNet-110 350 $93.02$
Diet-SNN [30] ResNet-20 10 $92.54$
BRP-SNN [74] CNN 20 $57.08$
Tandem Learning [71] CifarNet 8 $90.98$
SAC [75] ResNet-18 6 $93.74$
@@
CIFAR100 S-ResNet [69] ResNet-110 350 $70.62$
Diet-SNN [30] ResNet-20 5 $64.07$
@@ IV-C Results on CIFAR10-DVS
Neuromorphic datasets generally possess higher noise levels in comparison to static datasets, which heightens the risk of overfitting in well-optimized SNNs. Among all mainstream neuromorphic datasets, CIFAR10-DVS is recognized as one of the most challenging, presenting approximately 900 training samples for each label. Recent literature suggests an inclination towards complex architectures to handle this dataset effectively, albeit this strategy carries an inherent risk of overfitting without correspondingly enhancing the model’s accuracy. Therefore, a pressing need exists for novel methodologies that can reduce the influence of noise and tackle the difficulties inherent in working with CIFAR10-DVS.
Similar to the CIFAR10/100 datasets, we employ the spiking MLP-SPE-T network, pre-trained on the ImageNet-1K dataset, as the pre-trained model for the CIFAR10-DVS dataset. We reset the final classification output layer of the network, and subsequently fine-tune it on the resized $224\times 224$ CIFAR10-DVS dataset. Our fine-tuning strategy employs a cosine decay learning rate, starting with an initial value of 0.1 and progressively reducing to zero over 100 epochs. SGD, with a momentum of 0.9, is used as the optimizer.
@@
Table III: Comparison on CIFAR10-DVS (T denotes simulation length)